# Discrete Max-Focusing

### v2

@@ 5.3 Accessible Wedges and Accessible Throats
    $f\supset a$, ð~$\tilde{\eth}f=\tilde{\eth}a$; [^11] We are mainly interested here in the case where $a$ shares no boundary with conformal infinity, ð~$\tilde{\eth}a=\varnothing$. The stated condition allows us to be more general and also consider $a$ whose conformal edge is a subset of the conformal boundary in asymptotically AdS spacetimes. This is important in order to recover the traditionally notion of entanglement wedges of AdS boundary regions as a special case Bousso:2023sya . The condition ð~$\tilde{\eth}f=\tilde{\eth}a$ is not suitable when the spacetime has a different asymptotic structure. We thank S. Kaya for sharing ongoing work that will supply a more general condition, as well as an alternate definition of $e_{\rm min}$ as a complement of a suitable $e_{\rm max}$.
@@
    $f$ admits a Cauchy slice $\Sigma$ such that $\Sigma\supset\eth a$ and such that for any wedge $h\subsetneq f$ with $a\subset h$, ð$\eth h\subset\Sigma$, and ð$\eth h\setminus\eth f$ compact in $M$,
@@
Let $a$ be a wedge. A wedge $k$ is called future-marginally accessible from $a$ if $k$ is accessible from $a$ and $k$ is future-marginal on ð$\eth k\setminus\eth a$. A wedge $k$ is called past-marginally accessible from $a$ if $k$ is accessible from $a$ and $k$ is past-marginal on ð$\eth k\setminus\eth a$. (See the previous definition and Def. 49.)
@@
Let $a$ be a wedge. A wedge $k$ is called a throat accessible from $a$ if $k$ is past- and future-marginally accessible from $a$. (See the previous definition and Def. 50.) Equivalently, a throat $k$ for $a$ is called accessible if $k$ is accessible from $a$.
@@
In the limit where $a$ is an asymptotic region near the boundary of Anti-de Sitter space and where we approximate the conditional max entropy by a difference of von Neumann entropies, the the above definition reduces to the notion of an “outer-minimizing” wedge for a given asymptotic boundary subregion. The word “outer” here refers to the side of the throat quantum extremal surface ð$\eth k$ on which property III above is met. Here we take the wedge $k$ as the more fundamental object, from which ð$\eth k$ arises as its edge; so the terminology “outer” is obsolete.
@@ 6 Generalized Entanglement Wedges and Their Properties
Given a wedge $a$, let $F(a)$ be the set of all wedges accessible from $a$ (see Def. 53). The max entanglement wedge of $a$, $e_{\rm max}(a)$, is their wedge union:
@@
We must show that $e_{\rm max}(a)\in F(a)$, i.e., that $e_{\rm max}(a)$ satisfies properties I–III listed in Def. 53.
@@
Let $h\supset a$, ð$\eth h\subset\Sigma_{3}$. By property III of $\Sigma_{1}$ and $\Sigma_{2}$,
@@
By Discrete Max-Focusing (Conj. 32),
@@
All antinormal portions of ð$\eth e_{\rm max}(a)$ are throats. In particular, $e_{\rm max}(a)$ is an accessible throat.
@@
We will show that $e_{\rm max}(a)$ is future-marginal where it is antinormal; the proof that it is past-marginal is analogous. Suppose for contradiction that $e_{\rm max}(a)$ is antinormal but not future-marginal at $p\in\eth e_{\rm max}(a)$. Then for any open set $O$ containing $p$ there exists a future-nonexpanding past outward deformation $f_{1}$ of $e_{\rm max}(a)$ such that ð$\eth f_{1}\in\eth e_{\rm max}(a)\cup[O\cap H^{-}(e_{\rm max}(a)^{\prime})]$. Since $e_{\rm max}(a)$ is antinormal on the open set ð$\eth e_{\rm max}(a)\setminus\eth a$, $O$ can be chosen such that $e_{\rm max}(a)$ is antinormal on $O\cap\eth e_{\rm max}(a)$. Then by Discrete Max-Focusing, $f_{1}$ is also past-nonexpanding. Hence $f_{1}$ satisfies properties I and II of Def. 53. We will now show that $f_{1}$ satisfies property III with Cauchy surface $\Sigma_{1}=\Sigma_{\rm max}\cup H^{-}(e_{\rm max}(a)^{\prime})\cap f_{1}$, where $\Sigma_{\rm max}$ is the Cauchy surface of $e_{\rm max}(a)$ on which property III is satisfied.
Let $h\subset f_{1}$ with ð$\eth h\subset\Sigma_{1}$, $h\supset a$. By Property III of $e_{\rm max}(a)$,
@@
By Discrete Max-Focusing, Conj. 32,
@@
The second part of the theorem follows since $e_{\rm max}(a)$ is antinormal on ð$\eth e_{\rm max}(a)\setminus\eth a$, by Theorem 60. ∎
@@
    $g^{\prime}$ admits a Cauchy slice $\Sigma^{\prime}$ such that for any wedge $h\supsetneq g$ with ð$\eth h\subset\Sigma^{\prime}$ and ð$\eth h\setminus\eth g$ compact in $M$,
@@
We must show that $e_{\rm min}(a)$ satisfies properties i–iii listed in Def. 62.
@@
Property ii: The intersection of two normal wedges is normal by Lemma 47. Hence $e_{\rm min}(a)$ is normal by Def. 62.
@@
Let $h\supset g_{3}$, ð$\eth h\subset\Sigma^{\prime}_{3}$. By property iii of $\Sigma^{\prime}_{1}$ and $\Sigma^{\prime}_{2}$,
@@
By Discrete Max-Focusing (Conj. 32),
@@
Theorem 64.
$e_{\rm min}(a)^{\prime}$ is future- and/or past-marginal on ð$\eth e_{\rm min}(a)\setminus\eth a$. Specifically,
  • $e_{\rm min}(a)^{\prime}$ is future-marginal on ð$\eth e_{\rm min}(a)\cap H^{+}(a^{\prime})$.
  • $e_{\rm min}(a)^{\prime}$ is past-marginal on ð$\eth e_{\rm min}(a)\cap H^{-}(a^{\prime})$.
  • $e_{\rm min}(a)^{\prime}$ is a throat on ð$\eth e_{\rm min}(a)\cap a^{\prime}$.
Proof.
Let $p\in\eth e_{\rm min}(a)\cap H^{+}(a^{\prime})$. By Theorem 63, $e_{\rm min}(a)$ is normal at $p$ and so $e_{\rm min}(a)^{\prime}$ is future-nonexpanding at $p$. Suppose for contradiction that $e_{\rm min}(a)^{\prime}$ is not future-marginal at $p$. By arguments analogous to the proof of Theorem 61, for any open set $O$ containing $p$ there exists a future-nonexpanding past outward deformation $g_{1}^{\prime}$ of $e_{\rm min}(a)^{\prime}$ such that ð$\eth g_{1}\in\eth e_{\rm min}(a)^{\prime}\cup[O\cap H^{-}(e_{\rm min}(a))]$ whose complement $g_{1}$ still satisfies properties i-iii of Def. 62. Since $g_{1}$ is a proper subset of $e_{\rm min}(a)$, this contradicts the definition of $e_{\rm min}(a)$. The time-reversed argument applies to $O\subset\eth e_{\rm min}(a)\cap H^{-}(a^{\prime})$.
For $O\subset\eth e_{\rm min}(a)\cap a^{\prime}$, we similarly deform ð$\eth e_{\rm min}(a)$ inward (i.e., towards $a$) within $O$ along a past null geodesic to show that it is future-marginal, and inward along a future null geodesic to show that it is past-marginal. ∎
@@
00 0 $>H_{\rm max,gen}[(e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}|e_{\rm min}]$ (71)
@@
Special case 2: Suppose that $\Sigma_{\rm max}\subset J(\Sigma_{\rm min}^{\prime})$, i.e., $\Sigma_{\rm max}$ and $\Sigma_{\rm min}^{\prime}$ are everywhere causally separated. By Discrete Max-Focusing (Conj. 32),
@@
00 0 $>H_{\rm max,gen}[(e_{\rm max}{}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}|e_{\rm max}]+H_{\rm max,gen}[e_{\rm max}|e_{\rm min}\cap\Sigma_{\rm max}]$ (77)
@@
By Discrete Max-Focusing (Conj. 32),
@@
00 0 $>H_{\rm max,gen}[(e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}\cap(e_{\rm max}^{\prime}\cap e_{\rm min})^{\prime}|e_{\rm max}]+H_{\rm max,gen}[e_{\rm max}|e_{\rm min}\cap\Sigma_{\rm max}]$
@@
00 0 $>-H_{\rm max,gen}[(e_{\rm min}\cap\Sigma_{\rm max})^{\prime}\cap(e_{\rm min}\cap e_{\rm max}^{\prime})^{\prime}|e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime}]$
@@
If $a\subset b$ and ð$\eth e_{\rm min}(b)\setminus e_{\rm min}(a)$ is compact, then
@@
We will show that $g$ satisfies properties i-iii listed in Def. 62. This contradicts the definition of $e_{\rm min}(b)$ unless $g=e_{\rm min}(b)$, which is equivalent to the conclusion.
@@
This is a Cauchy slice of $g^{\prime}$. Let $h\supset g$ with ð$\eth h\subset\Sigma^{\prime}$. By property iii of $\Sigma_{\rm min}^{\prime}(b)$ and property III of $\Sigma_{\rm max}(a)$
@@
By Discrete Max-Focusing (Conj. 32),
@@
Note that ð$\eth x$ is nowhere to the past or future of ð$\eth e(bc)$. Therefore, there exists a single Cauchy slice that contains the edges of $x$, $e(bc)$, $x\cap e(bc)$, and $x\Cup e(bc)$. We note that $x\cap e(bc)=e(ab)\cap e(bc)$ and that $x\Cup e(bc)\subset e(ab)\Cup e(bc)$. By Theorem 66,
@@
Although $x$ need not be antinormal, Eq. (99) ensures that [ð$[\eth x\setminus\eth e(ab)]\cap\eth(x\Cup e(bc))=\varnothing$. Arguments analogous to the proof of Corollary 48 then imply that $x\Cup e(bc)$ is antinormal, except where its edge coincides with ð$\eth[e(abc)]$ and hence with $\Sigma[e(abc)]$. By Discrete Max-Focusing, property III of $e(abc)$, and the chain rule Eq. (14),

### v3

@@ 5.3 Accessible Wedges and Accessible Throats
    $f\supset a$, $\tilde{\eth}f=\tilde{\eth}a$; [^11] This condition will be updated in future work, for generality. Here as in Ref. Bousso:2023sya , we are mainly interested in the cases where $a$ (even if it shares no boundary with conformal infinity) is a wedge in an asymptotically AdS spacetime. The condition $\tilde{\eth}f=\tilde{\eth}a$ is not suitable when the spacetime has a different asymptotic structure. A more general condition will be supplied in forthcoming work BoussoKaya , alongside an alternate definition of $e_{\rm min}$ as a complement of a suitable $e_{\rm max}$.
@@
    $f$ admits a Cauchy slice $\Sigma$ such that $\Sigma\supset\eth a$ and such that for any wedge $h\subsetneq f$ with $a\subset h$, $\eth h\subset\Sigma$, and $\eth h\setminus\eth f$ compact in $M$,
@@
Let $a$ be a wedge. A wedge $k$ is called future-marginally accessible from $a$ if
  1. I.
    $k$ is accessible from $a$;
  2. II.
    $k$ is future-noncontracting.
Past-marginally accessible is defined analogously.
@@
Let $a$ be a wedge. A wedge $k$ is called a throat accessible from $a$ if $k$ is past- and future-marginally accessible from $a$. (See the previous definition and Def. 51.)
@@
In the limit where $a$ is an asymptotic region near the boundary of Anti-de Sitter space and where we approximate the conditional max entropy by a difference of von Neumann entropies, the the above definition reduces to the notion of an “outer-minimizing” wedge for a given asymptotic boundary subregion. The word “outer” here refers to the side of the throat quantum extremal surface $\eth k$ on which property III above is met. Here we take the wedge $k$ as the more fundamental object, from which $\eth k$ arises as its edge; so the terminology “outer” is obsolete.
@@ 6 Generalized Entanglement Wedges and Their Properties
Given a wedge $a$, let $F(a)$ be the set of all wedges accessible from $a$ (see Def. 54). The max entanglement wedge of $a$, $e_{\rm max}(a)$, is their wedge union:
@@
We must show that $e_{\rm max}(a)\in F(a)$, i.e., that $e_{\rm max}(a)$ satisfies properties I–III listed in Def. 54.
@@
Let $h\supset a$, $\eth h\subset\Sigma_{3}$. By property III of $\Sigma_{1}$ and $\Sigma_{2}$,
@@
By Discrete Max-Focusing (Conj. 33),
@@
$e_{\rm max}(a)$ is an accessible throat.
@@
We will show that $e_{\rm max}(a)$ is future-marginally accessible; the proof that it is past-marginally accessible is analogous. Suppose for contradiction that $e_{\rm max}(a)$ is antinormal on $\eth e_{\rm max}(a)\setminus\eth a$ but not future-marginal. Then for any open set $O\supset\eth e_{\rm max}(a)$, there exists a future-nonexpanding past-directed outward deformation $f_{1}$ of $e_{\rm max}(a)$ such that $\eth f_{1}\subset\eth e_{\rm max}(a)\cup[O\cap H^{-}(e_{\rm max}(a)^{\prime})]$. Since $e_{\rm max}(a)$ is antinormal on the open set $\eth e_{\rm max}(a)\setminus\eth a$, $O$ can be chosen such that $e_{\rm max}(a)$ is antinormal on $O\cap\eth e_{\rm max}(a)$. Then by Discrete Max-Focusing, $f_{1}$ is also past-nonexpanding. Hence $f_{1}$ satisfies properties I and II of Def. 54. We will now show that $f_{1}$ satisfies property III with Cauchy surface $\Sigma_{1}=\Sigma_{\rm max}\cup H^{-}(e_{\rm max}(a)^{\prime})\cap f_{1}$, where $\Sigma_{\rm max}$ is the Cauchy surface of $e_{\rm max}(a)$ on which property III is satisfied.
Let $h\subset f_{1}$ with $\eth h\subset\Sigma_{1}$, $h\supset a$. By Property III of $e_{\rm max}(a)$,
@@
By Discrete Max-Focusing, Conj. 33,
@@
The second part of the theorem follows since $e_{\rm max}(a)$ is antinormal on $\eth e_{\rm max}(a)\setminus\eth a$, by Theorem 61. ∎
@@
    $g^{\prime}$ admits a Cauchy slice $\Sigma^{\prime}$ such that for any wedge $h\supsetneq g$ with $\eth h\subset\Sigma^{\prime}$ and $\eth h\setminus\eth g$ compact in $M$,
@@
We must show that $e_{\rm min}(a)$ satisfies properties i–iii listed in Def. 63.
@@
Property ii: The intersection of two normal wedges is normal by Lemma 48. Hence $e_{\rm min}(a)$ is normal by Def. 63.
@@
Let $h\supset g_{3}$, $\eth h\subset\Sigma^{\prime}_{3}$. By property iii of $\Sigma^{\prime}_{1}$ and $\Sigma^{\prime}_{2}$,
@@
By Discrete Max-Focusing (Conj. 33),
@@
$0$ $>H_{\rm max,gen}[(e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}|e_{\rm min}]$ (71)
@@
Special case 2: Suppose that $\Sigma_{\rm max}\subset J(\Sigma_{\rm min}^{\prime})$, i.e., $\Sigma_{\rm max}$ and $\Sigma_{\rm min}^{\prime}$ are everywhere causally separated. By Discrete Max-Focusing (Conj. 33),
@@
$0$ $>H_{\rm max,gen}[(e_{\rm max}{}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}|e_{\rm max}]+H_{\rm max,gen}[e_{\rm max}|e_{\rm min}\cap\Sigma_{\rm max}]$ (77)
@@
By Discrete Max-Focusing (Conj. 33),
@@
$0$ $>H_{\rm max,gen}[(e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime})^{\prime}\cap(e_{\rm max}^{\prime}\cap e_{\rm min})^{\prime}|e_{\rm max}]+H_{\rm max,gen}[e_{\rm max}|e_{\rm min}\cap\Sigma_{\rm max}]$
@@
$0$ $>-H_{\rm max,gen}[(e_{\rm min}\cap\Sigma_{\rm max})^{\prime}\cap(e_{\rm min}\cap e_{\rm max}^{\prime})^{\prime}|e_{\rm max}^{\prime}\cap\Sigma_{\rm min}^{\prime}]$
@@
If $a\subset b$ and $\eth e_{\rm min}(b)\setminus e_{\rm min}(a)$ is compact, then
@@
We will show that $g$ satisfies properties i-iii listed in Def. 63. This contradicts the definition of $e_{\rm min}(b)$ unless $g=e_{\rm min}(b)$, which is equivalent to the conclusion.
@@
This is a Cauchy slice of $g^{\prime}$. Let $h\supset g$ with $\eth h\subset\Sigma^{\prime}$. By property iii of $\Sigma_{\rm min}^{\prime}(b)$ and property III of $\Sigma_{\rm max}(a)$
@@
By Discrete Max-Focusing (Conj. 33),
@@
Note that $\eth x$ is nowhere to the past or future of $\eth e(bc)$. Therefore, there exists a single Cauchy slice that contains the edges of $x$, $e(bc)$, $x\cap e(bc)$, and $x\Cup e(bc)$. We note that $x\cap e(bc)=e(ab)\cap e(bc)$ and that $x\Cup e(bc)\subset e(ab)\Cup e(bc)$. By Theorem 66,
@@
Although $x$ need not be antinormal, Eq. (99) ensures that $\eth x\setminus\eth e(ab)]\cap\eth(x\Cup e(bc))=\varnothing$. Arguments analogous to the proof of Corollary 49 then imply that $x\Cup e(bc)$ is antinormal, except where its edge coincides with $\eth[e(abc)]$ and hence with $\Sigma[e(abc)]$. By Discrete Max-Focusing, property III of $e(abc)$, and the chain rule Eq. (14),